OpenStudy (anonymous):

what is the factor? x^2-13x+40

OpenStudy (er.mohd.amir):

take 40 make two factor of 40 such that sum is 13 8*5=40 x^2-(5+8)x+40 =x^2-5x-8x+40 foctor it

OpenStudy (anonymous):

is it (x-5) (x-8)